OpenStudy (fibonaccichick666):

Let \(f:[a,b]\rightarrow R\) be a continuous function. Let \(\epsilon >0\) be a constant. For \(x \in[a+\epsilon,b-\epsilon]\), define \(g(x):=\frac{1}{2\epsilon} \int_{x-\epsilon}^{x+\epsilon} f\) a. Show that g is differentiable and find the derivative. b. Let f be differentiable and fix \(x\in(a,b)\)(let epsilon be small enough

OpenStudy (anonymous):

This an application of the fundamental theorem of calculus http://en.wikipedia.org/wiki/Fundamental_theorem_of_calculus

OpenStudy (anonymous):

That would seem to be my initial idea.

OpenStudy (fibonaccichick666):

I understand the FTC, but I fail to see how it is used to prove that f is differentiable

OpenStudy (fibonaccichick666):

I think we could use that for part b, because the integral will go to zero while 1/2epsilon will go to infinity, but... yea

OpenStudy (fibonaccichick666):

should we input g into the limit definition of derivative?

OpenStudy (anonymous):

The derivative is \[ g'(x)=\frac 1 {2 \epsilon} (f(x+\epsilon)- f(x+\epsilon)) \]

OpenStudy (anonymous):

The derivative is \[ g'(x)=\frac 1 {2 \epsilon} (f(x+\epsilon)- f(x-\epsilon)) \]

OpenStudy (fibonaccichick666):

how did you get that?

OpenStudy (fibonaccichick666):

since we don't know the variable of integration

OpenStudy (fibonaccichick666):

ohohohohoh

OpenStudy (anonymous):

Variable of integration is some dummy variable like \(t\).

OpenStudy (fibonaccichick666):

ok, nvm, I see it now

OpenStudy (anonymous):

The only problem is the \(\epsilon\) part.

OpenStudy (fibonaccichick666):

that's just a constant in the first case

OpenStudy (anonymous):

Actually, since the epsilon simply shrink the interval, I suppose it isn't a problem at all.

OpenStudy (anonymous):

Apply the FTC for the interval \( ( x-\epsilon, \quad c)\) and \( (c , \quad x+\epsilon) \)

OpenStudy (anonymous):

where c is in between. \(\epsilon \) is a constant

OpenStudy (anonymous):

Okay, but I think we can formalize this better.

OpenStudy (anonymous):

I am sorry, I have to go to a meeting. I think, you should be able to continue this problem

OpenStudy (anonymous):

Since \(f\) is continuous on \([a,b]\), it is also integrable.

OpenStudy (anonymous):

Integrable on \((a,b)\). We have defined \(g(x)\) to have a domain of \((a,b)\) or less.

OpenStudy (fibonaccichick666):

ok, so would this be a cohesive argument: \[g(x):=\frac{1}{2\epsilon} \int_{x-\epsilon}^{x+\epsilon} f\]\[=\frac{1}{2\epsilon} [F(x+\epsilon)-F(x-\epsilon)]~~by~ the~FTC\]now since we know that F is differentiable since it is the antiderivative of f, we find that \[g'(x)=:=\frac{1}{2\epsilon} [f(x+\epsilon)-f(x-\epsilon)] \]

OpenStudy (fibonaccichick666):

is that enough?

OpenStudy (fibonaccichick666):

thanks for the help elias

OpenStudy (anonymous):

Differentiable has a definition, but in general if the derivative is continuous, then it is differentiable.

OpenStudy (fibonaccichick666):

so do I need to add that this shows continuity?

OpenStudy (anonymous):

Differentiable means that the limit: \[ \lim_{h\to 0}\frac{f(x+h)-f(x)}{h} \]exists.

OpenStudy (fibonaccichick666):

but without a function, I don't understand how I would show that here, because I don't know how to apply f(x+h)

OpenStudy (fibonaccichick666):

I guess I also don't see how things will cancel

OpenStudy (anonymous):

All you need is the fundamental theorem of calculus. You are not required to prove it and the function meets the conditions required to apply it. If the FTC was proven in class you are not required to reprove it from first principles.

OpenStudy (fibonaccichick666):

hmm ok, i guess it just doesn't seem like enough

OpenStudy (anonymous):

To use FTC, you have to prove that \(x+\epsilon\) and \(x-\epsilon\) are in \((a,b)\).

OpenStudy (anonymous):

It's possible that \(x+\epsilon = (b-\epsilon)+\epsilon = b\)

OpenStudy (anonymous):

Wait, I suppose \(x<b-\epsilon\), so we can say \(x+\epsilon < b\)

OpenStudy (anonymous):

Anyway, the FTC is enough, just carefully state why the conditions are met. In particular you need \(f\) to be continuous on the interval it is being integrated on. In this case it is.

OpenStudy (anonymous):

At the very least, we can say \(g\) is differentiable on \((a+\epsilon,b-\epsilon)\).

OpenStudy (fibonaccichick666):

hmmm.... so is what I said before enough?

OpenStudy (anonymous):

Yeah, I think so. I'm looking at this: http://en.wikipedia.org/wiki/Fundamental_theorem_of_calculus#First_part

OpenStudy (freckles):

fibo you could use that definition that wio mentioned to find the derivative of g assume F'=f \[g'(x)=\lim_{h \rightarrow 0}\frac{g(x+h)-g(x)}{h} \\ =\lim_{h \rightarrow 0}\frac{\frac{1}{2 \epsilon } \int\limits_{x+h-\epsilon}^{x+h+\epsilon}f dx-\frac{1}{2 \epsilon} \int\limits_{x- \epsilon}^{x+\epsilon} f dx }{h} \\ =\frac{1}{2\epsilon} \lim_{h \rightarrow 0}\frac{F(x+\epsilon+h)-F(x-\epsilon+h)-F(x+\epsilon)+F(x-\epsilon)}{h} \\ =\frac{1}{ 2 \epsilon } \lim_{h \rightarrow 0}(\frac{F(x+\epsilon+h)-F(x+\epsilon)}{h}-\frac{F(x-\epsilon+h)-F(x-\epsilon)}{h}) \\=\frac{1}{2 \epsilon}(F'(x+\epsilon)-F'(x-\epsilon)) \\ =\frac{1}{2 \epsilon} (f(x+\epsilon)-f(x-\epsilon))\] but i know you guys are way passed this point now and sorry for butting in so late

OpenStudy (fibonaccichick666):

ok, so for part b, does this work? since the smallest epsilon can possibly get is 0, we can look at the limit as epsilon goes to zero\[\lim_{\epsilon\rightarrow 0} g'(x)=\lim_{\epsilon\rightarrow 0} \frac{1}{2\epsilon}[f(x+\epsilon)-f(x-\epsilon)]=\]\[\ [\lim_{\epsilon\rightarrow 0} \frac{1}{2\epsilon} ]\lim_{\epsilon\rightarrow 0} f(x+\epsilon)-f(x-epsilon)=\lim_{\epsilon\rightarrow 0} \frac{1}{2\epsilon} *0=0 \]

OpenStudy (fibonaccichick666):

thanks freckles, I just didn't know how that would work. but I am able to see it now

OpenStudy (anonymous):

Let \(y = x-\epsilon\) for a moment. \[ g'(x) = \lim_{\epsilon\to0}\frac{f(y+2\epsilon)-f(y)}{2\epsilon} \]

OpenStudy (fibonaccichick666):

it was that second to last step that got me

OpenStudy (fibonaccichick666):

ok

OpenStudy (anonymous):

Let \(h=2\epsilon\).

OpenStudy (fibonaccichick666):

then we have \[g'(x) = \lim_{\epsilon\to0}\frac{f(y+h)-f(y)}{h}\]

OpenStudy (freckles):

\[f'(y) = \lim_{\epsilon\to0}\frac{f(y+2\epsilon)-f(y)}{2\epsilon}\] ?

OpenStudy (fibonaccichick666):

which is just the def of the derivative

OpenStudy (anonymous):

Maybe I should have still put: \[ \lim_{\epsilon\to 0}g'(x) \]

OpenStudy (freckles):

yes where \[\lim_{\epsilon \rightarrow 0}g'(x)=f'(x) \text{ right? }\]

OpenStudy (anonymous):

Yeah, the only part that is a bit tedious is we have \(\lim_{\epsilon\to0}\) instead of \(\lim_{2\epsilon\to0}\).

OpenStudy (fibonaccichick666):

2*0 is still zero though...?

OpenStudy (anonymous):

There is some composition of limits rule that would work for it though.

OpenStudy (fibonaccichick666):

so i take it my explanation isn't enough?

OpenStudy (anonymous):

It's good enough I think.

OpenStudy (anonymous):

Maybe you could say: \[ \lim_{\epsilon\to 0}2\epsilon = 0 \]

OpenStudy (fibonaccichick666):

ok, I can add that, but why would it be necessary? I'm not computing that limit to avoid the infinity

OpenStudy (anonymous):

There isn't any rule about what level of rigor is necessary.

OpenStudy (fibonaccichick666):

no, but the more the better in this class

OpenStudy (fibonaccichick666):